If Tedd can run 5 miles in 56 minutes how many miles would he run in 84 minutes? Can somebody help me set this word problem up? I have tried everything and can't get it right.
I also have another word problem that I am fumbled with.
There are 630 dishes that need washed. Jim can do it in 105 minutes. His friend can do it in 70 minutes. If they do it together how many minutes will it take to rinse 630 dishes?? Please help! Ty in advance. Judy

Answer by checkley75(3666)

5/56=x/84
56x=5*84
56x=420
x=420/56
7.5 miles in 84 minutes.